# EarthMarker: A Visual Prompt Learning Framework for Region-level and Point-level Remote Sensing Imagery Comprehension

### v1

@@ Abstract
Recent advances in visual prompting in the natural image area have allowed users to interact with artificial intelligence (AI) tools through various visual marks such as box, point, and free-form shapes. However, due to the significant difference between the natural and remote sensing (RS) images, existing visual prompting models face challenges in RS scenarios. Moreover, RS MLLMs mainly focus on interpreting image-level RS data and only support interaction with language instruction, restricting flexibility applications in the real world. To address those limitations, a novel visual prompting model named EarthMarker is proposed, which excels in image-level, region-level, and point-level RS imagery interpretation. Specifically, the visual prompts alongside images and text instruction input into the large language model (LLM), adapt models toward specific predictions and tasks. Subsequently, a sharing visual encoding method is introduced to refine multi-scale image features and visual prompt information uniformly. Furthermore, to endow the EarthMarker with versatile multi-granularity visual perception abilities, the cross-domain phased learning strategy is developed, and the disjoint parameters are optimized in a lightweight manner by leveraging both the natural and RS domain-specific knowledge. In addition, to tackle the lack of RS visual prompting data, a dataset named RSVP featuring multi-modal fine-grained visual prompting instruction is constructed. Extensive experiments are conducted to demonstrate the proposed EarthMarker’s competitive performance, representing a significant advance in multi-granularity RS imagery interpretation under the visual prompting learning framework. Our code and dataset are available at https://github.com/wivizhang/EarthMarker.
@@
In the RS field, there are a few works that limited works were devoted to region-level imagery interpretation. For example, RSVG [16] adopts language prompting to inquire and localize the specific object, but does not involve the visual prompting technique to realize various visual tasks. In addition, inspired by SAM, RSPrompter [17] introduces an automated prompts generation to develop interactive segmentation in RS imagery. Another representative work is EarthGPT [5], which shows the potential of region-level image comprehension by training on visual grounding datasets. However, EarthGPT only supports language interaction without visual prompts, thereby lacking flexibility. These limitations hinder the development of fine-grained spatial understanding and complex reasoning execution in the RS domain. It is clear that the visual prompts learning in the RS domain is still in its infancy.
To bridge this gap, a fine-grained MLLM named EarthMarker is proposed, leveraging visual prompting to extend the capability of MLLMs for region-level and point-level understanding in the RS domain for the first time. Based on the visual prompts learning, as illustrated in Fig. 1, EarthMarker excels at the multi-granularity interpretation of RS imagery across image, region, and point levels. Moreover, EarthMarker can complete a wide range of RS visual reasoning tasks, including scene classification, referring object classification, captioning, inter-relationship analysis, etc. Concretely, the visual prompts, i.e., bounding boxes and points, along with the RS images and the text instructions are provided as input to the LLM. Notably, the visual prompting marks are utilized to isolate specific areas and guide the model to interpret regional content in the entire RS image. Considering that the RS imagery is gathered from an overhead perspective by satellites, associated with large-scale variations and cluttered backgrounds, multi-resolution image input processing is necessary. Subsequently, unlike most existing nature scene visual prompting works, which routinely set different visual encoders and visual prompts encoders. In our method, a sharing visual encoding method is developed. Specifically, the visual prompt is processed to RGB images analogously, which shares the same visual encoder with the inputted image. This strategy is beneficial for consistent feature extraction and understanding the relationship between visual prompts regions and the holistic image, enhancing the performance of the model under visual prompts learning.
@@
Extensive tests are conducted on multi-type RS datasets to evaluate the performance of EarthMarker which is demonstrated to be superior to state-of-the-art (SOTA) specialist models, MLLMs, and visual prompting models in various RS visual tasks at different granularity. Specifically, for the zero-shot scene classification task, EarthMarker shows a significant improvement compared with other existing MLLMs. Notably, for referring object classification, EarthMarker achieves a Semantic Similarity (SS) score of 98.37 using bounding boxes as visual prompts and 95.96 using point prompts on DIOR-RSVG dataset[19]. Furthermore, for image and region captioning tasks, EarthMarker also far exceeds other MLLMs and visual prompting models on the NWPU-Captions[20] dataset. In summary, the experimental results demonstrate that EarthMarker exhibits exceptional performance across a variety of multi-granularity RS image comprehension tasks and excellent zero-shot reasoning capability.
@@
Superior performance on multi-granularity RS Visual Tasks. Extensive experiments are conducted to demonstrate EarthMarker’s competitive performance in multi-granularity RS visual interpretation tasks, compared with the SOTA specialist models, MLLMs, and visual prompting models. The tasks evaluated include scene classification, referring object classification, captioning, and inter-relationship analyses. Therefore, EarthMarker successfully explores the adaptation of the visual prompt learning framework in the RS domain, improving the performance of MLLM and representing a significant step in fine-grained RS imagery interpretation.
@@ III-A Model Architecture
In particular, the images and corresponding visual prompts share a visual encoding mechanism for feature sharing, enabling the visual encoders to better understand and associate the relationship between images and visual prompts. Specifically, the Mixture of Visual Experts (MoV) [36] is designed to encode the visual information. The MoV incorporates two visual encoders, DINOv2-ViT L/14 [37] and CLIP-ConvNeXt [38], which are pre-trained on distinct network architectures (ViT and CNN), thus offering complementary visual semantics. To refine the robust multi-scale visual features, the input images are downsampled to different resolutions denoted as and then fed into the MoV module to encode, respectively. Leveraging the strengths of various visual backbones, visual perception is enhanced and key details in images are refined. Subsequently, the encoded visual features are transformed to the same dimension and concatenated channel-wisely to obtain the integrated multi-scale feature maps represented as . This process can be formulated simply as
Notably, a key step to the encoder-sharing mechanism is the “Visual Prompt as Images”. Especially, the dimension () of the visual prompts is processed to the same dimension () with the images. Then, the transformed visual prompts also can be fed into MoV together with the images, the encoded visual prompts expressed as . Similarly, this process is written as
Subsequently, the modality alignment projection layer transforms the visual tokens into the language semantic space. Meanwhile, the text instructions are processed by the tokenizer module, which handles text tokenization and embedding, converting them into text embeddings . After obtaining the projected image tokens, visual prompts tokes, and text instructions embeddings, they are integrated into an entire multi-modal input sequence. LLM decoder takes the multi-modal inputs and generates the response sequence Y. Formulated as
We employ Llama 2, a transformer-based decoder-only LLM, as the LLM decoder.
@@ III-B Cross-domain Phased Training
RS Visual Prompting Tuning. The last stage focuses on accurately following user instructions and achieving complex region-level and point-level visual reasoning tasks. The MoV, alignment projection, and LLM are fixed. The LoRA method is adopted for tuning. We load the weights trained in the previous phase and continue training EarthMarker on RSVP-3M region-text and point-text parings, which contain the fine-grained referring object classification and referring brief caption data. Specifically, several learnable low-rank adapter matrices are inserted into Transformer layers of LLM. The adapted multi-head attention is denoted as , the output of the -th adapted Transformer attention is formulated as
In conclusion, the cross-domain phased training endows EarthMarker with various granular (e.g., image-level, point-level, and region-level) multimodal instruction capabilities in the RS domain. In the first multi-domain image-text alignment stage, the LLM is efficiently converted into an MLLM, which is capable of image-level understanding. Subsequently, by utilizing the nature scene referring datasets, EarthMarker is equipped with the fundamental spatial perception of images. This is beneficial for subsequent developments of referring ability in the RS domain. Furthermore, by leveraging the RS visual prompting datasets RSVP-3M, EarthMarker is endowed with image understanding at both the region level and point level. Notably, different field’s datasets are adopted for training, and enhancing open-vocabulary reasoning ability. Notably, during the whole training, our updatable parameters are disjoint, preventing interference between understanding images at different granularity and the capability to follow visual prompts.
@@ IV-A Data Conversion and Annotation from Public RS Datasets
Task | Raw Data | Samples |
@@
Referring Object Classification (Point) | Vaihingen | 45,104 |
Potsdam | 504,139 | |
@@
In RSVP-3M, each data item consists of visual prompts, user instructions, and images. The visual prompts in the user instructions or model answers are expressed as or . For the point-level data, for example, the user instruction which guides referring object classification is “Please identify the category of each marked point in the image”. The answer format is “”. In addition, for the region-level data, Take the scene of the airport as an example. The user instruction for airport region captioning is “Please provide the brief caption of each marked region in the image, and the corresponding answer format generated by the model is “ ”. The data structures of other visual tasks are similar to those explained above. Through the transformation and re-annotation based on public datasets, the visual prompting dataset RSVP-3M is effectively developed, featuring image-point-text and image-region-text pairings.
@@ IV-B GPT4V-assisted Visual Prompting Data Generation
The data generated using GPT-4V not only compensates for the lack of information in brief captions but also provides detailed descriptions that reveal the spatial and semantic relationships between different regions in the image. For example, in aerial imagery, it is feasible to identify the general category of the image and provide a simple description. Additionally, detailed descriptions, such as the spatial layout of tennis courts, basketball courts, and playgrounds, the relationships among these areas, and the activities of people on the playground, can be conducted. The RSVP-3M dataset, supplemented by public datasets and data generated by GPT-4V, covers a wide range of fine-grained visual reasoning tasks, enhancing the richness and diversity of the data. The structure of this subset’s data is as follows.
@@ V-A Implementation Details
The proposed EarthMarker adopts the cross-domain phased training strategy, and the parameters updated vary at different stages. In general, we train an off-the-shelf 13B language model Llama 2 and the visual encoder MoV is kept frozen during the training. In the first multi-domain image-text alignment stage, only the alignment projection layer is updated. Then, in the spatial perception tuning phase, only the attention layers of LLM are unfrozen. Furthermore, the trainable LoRA metrics are introduced in the last RS visual prompting tuning stage. We utilized AdamW optimizer[64] with weight decay = 0 and betas = (0.9, 0.95), the learning rate is set to 2e-5, and the total training stages are conducted on 8 NVIDIA A100 GPUs. For model evaluation, we select diverse multi-granularity visual tasks to assess the performance of EarthMarker. Image-level tasks include scene classification, image captioning, and region-level tasks contain referring object classification and region captioning.
Method | Publication Year | BLEU1 | BLEU2 | BLEU3 | BLEU4 | METEOR | ROUGE-L | CIDEr(0-5) | SPICE |
Expert Model | |||||||||
@@ V-B Scene Classification
We prompt the model with an image-level box to represent the entire image. The text instruction is “Please identify the object category of each marked region in the image.”. We calculate the zero-shot accuracy on the AID and UCMerced dataset. EarthMarker significantly outperforms other VLMs, with an accuracy of 86.52 on UCMerced and 77.97 on AID, as presented in Tab. II. In comparison, LLaVa-1.5 and Sphinx, due to the lack of RS domain knowledge, are inferior to the RS MLLM GeoChat and our EarthMarker. Compared to GeoChat, our EarthMarker achieved an accuracy improvement of 5.94 on AID and 2.09 on UCMerced.
Method | Publication Year | Format | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE | CIDER | SPICE |
@@
Method | Publication Year | Format | SS | SIOU |
MLLM | ||||
@@
Visual Prompting Model | ||||
@@ V-C Image Captioning
To evaluate the image captioning capabilities, we use the NWPU-Captions[20] dataset to assess and compare EarthMarker against other expert models in the supervised setting. Created by Northwestern Polytechnical University, The NWPU-Captions dataset includes 31,500 aerial images and 157,500 sentences for RS image description. Following the protocol of MLCA-Net, we employ BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and CIDErD as evaluation metrics. In the evaluation, we use the sentence “Please provide a brief caption of each marked region in the image.” as text instruction and a full-image box as the visual prompt. As shown in Tab. III, compared to other expert models, EarthMarker demonstrate improvements in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, and ROUGE-L by 7.4, 8.2, 8.8, 6.5, 3.5, and 9.9, respectively, and a 36.5 improvement in CIDErD.
@@ V-D Referring Object Classification
As the results shown in Tab. V, EarthMarker achieves 95.96 in SS and 93.49 in S-IoU using point-level visual prompts, and 98.37 in SS and 97.24 in S-IoU based on box-level visual prompts on the DIOR-RSVG dataset. Both sets of results significantly outperform the SOTA method. Furthermore, EarthMarker surpasses the previous SOTA model EarthGPT by 3.73 in SS and 7.08 in S-IoU on the DIOR-RSVG dataset, demonstrating its robust capability in fine-grained box-level classification. Additionally, in Fig. 3, there are considerable differences in predictions of regional object categories by EarthMarker and other MLLMs, as well as visual prompting models. It is evident that when faced with complex geographical scenes and blurry tiny objects, EarthMarker’s predictions are significantly accurate.

### v2

# EarthMarker: Visual Prompt Learning for Region-level and Point-level Remote Sensing Imagery Comprehension

@@ Abstract
Recent advances in visual prompting in the natural image area have allowed users to interact with artificial intelligence (AI) tools through various visual marks such as box, point, and free-form shapes. However, due to the significant difference between the natural and remote sensing (RS) images, existing visual prompting models face challenges in RS scenarios. Moreover, RS MLLMs mainly focus on interpreting image-level RS data and only support interaction with language instruction, restricting flexibility applications in the real world. To address those limitations, the first visual prompting model named EarthMarker is proposed, which excels in image-level, region-level, and point-level RS imagery interpretation. Specifically, the visual prompts alongside images and text instruction input into the large language model (LLM), adapt models toward specific predictions and tasks. Subsequently, a sharing visual encoding method is introduced to refine multi-scale image features and visual prompt information uniformly. Furthermore, to endow the EarthMarker with versatile multi-granularity visual perception abilities, the cross-domain phased learning strategy is developed, and the disjoint parameters are optimized in a lightweight manner by leveraging both the natural and RS domain-specific knowledge. In addition, to tackle the lack of RS visual prompting data, a dataset named RSVP featuring multi-modal fine-grained visual prompting instruction is constructed. Extensive experiments are conducted to demonstrate the proposed EarthMarker’s competitive performance, representing a significant advance in multi-granularity RS imagery interpretation under the visual prompting learning framework. Our code and dataset are available at https://github.com/wivizhang/EarthMarker.
@@
In the RS field, there are limited works devoted to region-level imagery interpretation. For example, RSVG [16] adopts language prompting to inquire and localize the specific object, but does not involve the visual prompting technique to realize various visual tasks. In addition, inspired by SAM, RSPrompter [17] introduces an automated prompts generation to develop interactive segmentation in RS imagery. Another representative work is EarthGPT [5], which shows the potential of region-level image comprehension by training on visual grounding datasets. However, EarthGPT only supports language interaction without visual prompts, thereby lacking flexibility. These limitations hinder the development of fine-grained spatial understanding and complex reasoning execution in the RS domain. It is clear that the visual prompts learning in the RS domain is still in its infancy.
To bridge this gap, a fine-grained MLLM named EarthMarker is proposed, leveraging visual prompting to extend the capability of MLLMs for region-level and point-level understanding in the RS domain for the first time. Based on the visual prompts learning, as illustrated in Fig. 1, EarthMarker excels at the multi-granularity interpretation of RS imagery across image, region, and point levels. Moreover, EarthMarker can complete a wide range of RS visual reasoning tasks, including scene classification, referring object classification, captioning, inter-relationship analyses, etc. Concretely, the visual prompts, i.e., bounding boxes and points, along with the RS images and the text instructions are provided as input to the LLM. Notably, the visual prompting marks are utilized to isolate specific areas and guide the model to interpret regional content in the entire RS image. Considering that the RS imagery is gathered from an overhead perspective by satellites, associated with large-scale variations and cluttered backgrounds, multi-resolution image input processing is necessary. Subsequently, unlike most existing nature scene visual prompting works, which routinely set different visual encoders and visual prompts encoders. In our method, a sharing visual encoding method is developed. Specifically, the visual prompt is processed to RGB images analogously, which shares the same visual encoder with the inputted image. This strategy is beneficial for consistent feature extraction and understanding the relationship between visual prompts regions and the holistic image, enhancing the performance of the model under visual prompts learning.
@@
Extensive experiments are conducted on multi-type RS datasets to evaluate the performance of EarthMarker which is demonstrated to be superior to state-of-the-art (SOTA) specialist models, MLLMs, and visual prompting models in various RS visual tasks at different granularity. Specifically, for the zero-shot scene classification task, EarthMarker shows a significant improvement compared with other existing MLLMs. Notably, for referring object classification, EarthMarker achieves a Semantic Similarity (SS) score of 98.37 using bounding boxes as visual prompts and 95.96 using point prompts on DIOR-RSVG dataset[19]. Furthermore, for image and region captioning tasks, EarthMarker also far exceeds other MLLMs and visual prompting models. In summary, the experimental results demonstrate that EarthMarker exhibits exceptional performance across a variety of multi-granularity RS image comprehension tasks and excellent zero-shot reasoning capability.
@@
Superior Performance on Multi-granularity RS Visual Tasks. Extensive experiments are conducted to demonstrate EarthMarker’s competitive performance in multi-granularity RS visual interpretation tasks, compared with the SOTA specialist models, MLLMs, and visual prompting models. The tasks evaluated include scene classification, referring object classification, captioning, and inter-relationship analyses. Therefore, EarthMarker successfully explores the adaptation of the visual prompt learning in the RS domain, improving the performance of MLLM and representing a significant step in fine-grained RS imagery interpretation.
@@ III-A Model Architecture
In particular, the images and corresponding visual prompts share a visual encoding mechanism for feature sharing, enabling the visual encoders to better understand and associate the relationship between images and visual prompts. Specifically, the Mixture of Visual Experts (MoV) [35] is designed to encode the visual information. The MoV incorporates two visual encoders, DINOv2-ViT L/14 [36] and CLIP-ConvNeXt [37], which are pre-trained on distinct network architectures (ViT and CNN), thus offering complementary visual semantics. To refine the robust multi-scale visual features, the input images are downsampled to different resolutions denoted as and then respectively fed into the MoV module to encode. Leveraging the strengths of various visual backbones, visual perception is enhanced and key details in images are refined. Subsequently, the encoded visual features are transformed to the same dimension and concatenated channel-wisely to obtain the integrated multi-scale feature maps represented as . This process can be formulated as
Notably, a key step to the encoder-sharing mechanism is the “Visual Prompt as Images”. Especially, the dimension () of the visual prompts is processed to the same dimension () with the images. Then, the transformed visual prompts P also can be fed into MoV together with the images, the encoded visual prompts expressed as . Similarly, this process is written as
Subsequently, the modality alignment projection layer transforms the visual tokens into the language semantic space. Meanwhile, the text instructions are processed by the tokenizer module, which handles text tokenization and embedding, converting them into text embeddings . After obtaining the projected image tokens, visual prompts tokes, and text instructions embeddings, they are integrated into an entire multi-modal input sequence. LLM decoder takes the multi-modal inputs and generates the response sequence Y, which can be formulated as
It should be noted that we employ Llama 2, a transformer-based decoder-only LLM, as the LLM decoder.
@@ III-B Cross-domain Phased Training
RS Visual Prompting Tuning. The last stage focuses on accurately following user instructions and achieving complex region-level and point-level visual reasoning tasks. The MoV, alignment projection, and LLM are fixed. The LoRA method is adopted for tuning. We load the weights trained in the previous phase and continue training EarthMarker on RSVP-3M region-text and point-text parings, which contain the fine-grained referring object classification and referring brief caption data. Specifically, several learnable low-rank adapter matrices , , and are inserted into Transformer layers of LLM. The is the number of attention heads. The adapted multi-head attention is denoted as , thus the output of the -th adapted Transformer attention is formulated as
In conclusion, the cross-domain phased training endows EarthMarker with various granular (e.g., image-level, point-level, and region-level) multimodal instruction capabilities in the RS domain. In the first multi-domain image-text alignment stage, the LLM is efficiently converted into an MLLM, which is capable of image-level understanding. Subsequently, by utilizing the nature scene referring datasets, EarthMarker is equipped with the fundamental spatial perception of images. This is beneficial for subsequent developments of referring ability in the RS domain. Furthermore, by leveraging the RS visual prompting datasets RSVP-3M, EarthMarker is endowed with image understanding at both the region level and point level. Notably, different field’s datasets are adopted for training, and enhancing open-vocabulary reasoning ability. It should be emphasized that during the whole training, our updatable parameters are disjoint, preventing interference between understanding images at different granularity and the capability to follow visual prompts.
@@ IV-A Data Conversion and Annotation from Public RS Datasets
Tasks | Raw Data | Samples |
@@
Referring Object Classification (Point) | Vaihingen [58] | 45,104 |
Potsdam [58] | 504,139 | |
@@
In RSVP-3M, each data item consists of visual prompts, user instructions, and image. The visual prompts in the user instructions or model answers are expressed as or . For the point-level data, for example, the user instruction which guides referring object classification is “Please identify the category of each marked point in the image”. The answer format is “‘points’”. In addition, for the region-level data, take the scene of the airport as an example. The user instruction for airport region captioning is “Please provide the brief caption of each marked region in the image, and the corresponding answer format generated by the model is “ ‘bbox’”. The data structures of other visual tasks are similar to those explained above. Through the transformation and re-annotation based on public datasets, the visual prompting dataset RSVP-3M is effectively developed, featuring image-point-text and image-region-text pairings.
@@ IV-B GPT4V-assisted Visual Prompting Data Generation
The data generated using GPT-4V not only compensates for the lack of information in brief captions but also provides detailed descriptions that reveal the spatial and semantic relationships between different regions in the image. For example, in aerial imagery, it is feasible to identify the general category of the image and provide a simple description. Additionally, detailed descriptions, such as the spatial layout of tennis courts, basketball courts, playgrounds, the relationships among these areas, and the activities of people on the playground, can be conducted. The RSVP-3M dataset, supplemented by public datasets and data generated by GPT-4V, covers a wide range of fine-grained visual reasoning tasks, enhancing the richness and diversity of the data.
@@ V-A Implementation Details
The proposed EarthMarker adopts the cross-domain phased training strategy, and the parameters updated vary at different stages. In general, we train an off-the-shelf 13B language model Llama 2 and the visual encoder MoV is kept frozen during the training. In the first multi-domain image-text alignment stage, only the alignment projection layer is updated. Then, in the spatial perception tuning phase, only the attention layers of LLM are unfrozen. Furthermore, the trainable LoRA metrics are introduced in the last RS visual prompting tuning stage. We utilized AdamW optimizer[65] with weight decay = 0 and betas = (0.9, 0.95), the learning rate is set to 2e-5, and the total training stages are conducted on 8 NVIDIA A100 GPUs. For model evaluation, we select diverse multi-granularity visual tasks to assess the performance of EarthMarker. Image-level tasks include scene classification and image captioning, while region-level tasks contain referring object classification, region captioning, and relationship analyses.
Methods | Publication Year | BLEU1 | BLEU2 | BLEU3 | BLEU4 | METEOR | ROUGE-L | CIDEr(0-5) | SPICE |
Expert Models | |||||||||
@@ V-B Scene Classification
We prompt the model with an image-level box to represent the entire image. The text instruction is “Please identify the object category of each marked region in the image”. We calculate the zero-shot accuracy on the AID and UCMerced dataset. EarthMarker significantly outperforms other VLMs, with an accuracy of 86.52 on UCMerced and 77.97 on AID, as presented in Tab. II. In comparison, LLaVa-1.5 and Sphinx, due to the lack of RS domain knowledge, are inferior to the RS MLLM GeoChat and our EarthMarker. Compared to GeoChat, our EarthMarker achieved an accuracy improvement of 5.94 on AID and 2.09 on UCMerced. Owing to the multi-domain image-text alignment training, EarthMarker is endowed with excellent holistic scene understanding ability on RS imagery.
Method | Publication Year | Formats | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE | CIDER | SPICE |
@@
Methods | Publication Year | Formats | SS | SIOU |
MLLMs | ||||
@@
Visual Prompting Models | ||||
@@ V-C Image Captioning
To evaluate the image captioning capabilities, we use the NWPU-Captions[41] dataset to assess and compare EarthMarker against other expert models in the supervised setting. Created by Northwestern Polytechnical University, the NWPU-Captions dataset includes 31,500 aerial images and 157,500 sentences for RS image description. Following the protocol of MLCA-Net, we employ BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and CIDErD as evaluation metrics. In the evaluation, we use the sentence “Please provide a brief caption of each marked region in the image.” as text instruction and a full-image box as the visual prompt. As shown in Tab. III, compared to other expert models, EarthMarker demonstrates improvements in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, and ROUGE-L by 7.4, 8.2, 8.8, 6.5, 3.5, and 9.9, respectively, and a 36.5 improvement in CIDErD. In summary, we dexterously annotate the entire image as a single region, and further using both natural and RS image training, EarthMarker achieves a high level of understanding of the overall image, outperforming expert models.
@@ V-D Referring Object Classification
As the results shown in Tab. V, EarthMarker achieves 95.96 in SS and 93.49 in S-IoU using point-level visual prompts, and 98.37 in SS and 97.24 in S-IoU based on box-level visual prompts on the DIOR-RSVG dataset. Both sets of results significantly outperform the SOTA method. Furthermore, EarthMarker surpasses the previous SOTA model EarthGPT by 3.73 in SS and 7.08 in S-IoU on the DIOR-RSVG dataset, demonstrating its superior capability in fine-grained box-level classification and the effectiveness of our designed visual prompting method. Additionally, in Fig. 3, there are considerable differences in predictions of regional object categories by EarthMarker and other MLLMs, as well as visual prompting models. It is evident that when faced with complex geographical scenes and blurry tiny objects, EarthMarker’s predictions are significantly accurate.